The accompnaying data were read from a graph that appeared in a recent study. The independent variable is SO2 deposition rate (mg/m2/day) and the dependent variable is steel weight loss (g/m2).

x 14 18 40 43 45 112
y 280 350 470 500 560 1200

Calculate the residual for when SO2 deposition rate is 40 mg/m2/day.
Answer #1
X Y X * Y X2 Residual = ( Y - Ŷ )
14 280 3920 196 268.2375685 11.76243153
18 350 6300 324 305.4838363 44.51616368
40 470 18800 1600 510.3383095 -40.3383095
43 500 21500 1849 538.2730104 -38.2730104
45 560 25200 2025 556.8961443 3.103855655
112 1200 134400 12544 1180.771131 19.22886908
Total 272 3360 210120 18538 3360 3.41061E-13

Equation of regression line is Ŷ = a + bX
b = ( n Σ(XY) - (ΣX* ΣY) ) / ( n Σ X2 - (ΣX)2 )
b = ( 6 * 210120 - 272 * 3360 ) / ( 6 * 18538 - ( 272 )2)
b = 9.3116

a =( ΣY - ( b * ΣX ) ) / n
a =( 3360 - ( 9.3116 * 272 ) ) / 6
a = 137.8756
Equation of regression line becomes Ŷ = 137.8756 + 9.3116 X

Residual for X = 40 is -40.3383095.